<div dir="ltr">Alexander, my response to your questions. -Richard<div><br></div><blockquote class="gmail_quote" style="margin:0px 0px 0px 0.8ex;border-left:1px solid rgb(204,204,204);padding-left:1ex">Comparing the two versions with and without the requirement of ADC/TDC<br>matching for the hodoscope, I only see a loss of events below 1% in<br>almost the full energy range. There is only one 100MeV-wide bin near<br>9.1GeV, where the loss is about 15%. Could this be a sign of a broken or<br>miscalibrated TDC?</blockquote><div><br></div><div>No, the DAQ was working ok in that sector. The loss of 15% is pretty much uniform across the entire TAGM. The reason the ratio looks like 1 in the low-energy end is because the missed tags are missing from both the numerator (tdc*adc) and denominator (adc) in that region, so the ratio looks close to 1. I will present my analysis of these data at the beamline meeting next Tuesday. </div><div><br></div><blockquote class="gmail_quote" style="margin:0px 0px 0px 0.8ex;border-left:1px solid rgb(204,204,204);padding-left:1ex"> Once the thresholds were fixed in run 71724, the loss goes down to 10%<br>in the higher-energy part of the microscope, while it completely<br>recovers for the low-energy part. What is the reason for this<br>discrepancy? New vs old fibers?</blockquote><div><br></div><div>Basically the ratio is close to 1 in the low-energy end of the TAGM because the low-amplitude pulses are missed by both the adc and tdc. What really surprised me is the additional tags that are recovered at the high-energy end where the new fibers are: by lowering the adc threshold way down, we are able to see an additional 15% of tags in the adc over what is seen by the tdc which just counts on the primary MIP peak. I believe that this 15% are coming from electrons that passed through the cladding of the fiber, instead of the core. Originally I assumed that light produced in the cladding would be totally lost because it is outside the total-internal-reflection capture cone of the fiber. Now I see that this assumption was wrong! Nevertheless, there is probably not enough light from these pulses to make good tags.</div><div><br></div><div>Let's discuss the "15% problem" at the beamline meeting next week.</div></div><br><div class="gmail_quote"><div dir="ltr" class="gmail_attr">On Fri, Apr 24, 2020 at 11:46 AM Alexander Austregesilo <<a href="mailto:aaustreg@jlab.org">aaustreg@jlab.org</a>> wrote:<br></div><blockquote class="gmail_quote" style="margin:0px 0px 0px 0.8ex;border-left:1px solid rgb(204,204,204);padding-left:1ex">*Message sent from a system outside of UConn.*<br>
<br>
<br>
Hi,<br>
<br>
I looked at the accidental corrected rho yields as a function of energy<br>
for the three test runs. You can see the plots on this wiki page:<br>
<br>
<a href="https://urldefense.proofpoint.com/v2/url?u=https-3A__nam10.safelinks.protection.outlook.com_-3Furl-3Dhttps-253A-252F-252Fhalldweb.jlab.org-252Fwiki-2Dprivate-252Findex.php-252F2020-5FTagM-5FThreshold-5FTest-26amp-3Bdata-3D02-257C01-257Crichard.t.jones-2540uconn.edu-257Ca352d79c11974c79fea808d7e866a12e-257C17f1a87e2a254eaab9df9d439034b080-257C0-257C0-257C637233399810094785-26amp-3Bsdata-3D4zYaxV2ZzXphRsiaR5eFzgX3ebxm-252BHN7owoyTIl-252F1lA-253D-26amp-3Breserved-3D0&d=DwMFaQ&c=CJqEzB1piLOyyvZjb8YUQw&r=-MwMW0sKroUIjy-Lh9rb3KzmWIgdcbbr9_jez8RLmto&m=iX1lXNpi6V_q0QPmEr59lqvAzNjcGLwUObwfy25WB3c&s=ANiDrG8v8JeYgM7CpEonES5NeseIQAkdvo2A7xZJYdU&e=" rel="noreferrer" target="_blank">https://nam10.safelinks.protection.outlook.com/?url=https%3A%2F%2Fhalldweb.jlab.org%2Fwiki-private%2Findex.php%2F2020_TagM_Threshold_Test&data=02%7C01%7Crichard.t.jones%40uconn.edu%7Ca352d79c11974c79fea808d7e866a12e%7C17f1a87e2a254eaab9df9d439034b080%7C0%7C0%7C637233399810094785&sdata=4zYaxV2ZzXphRsiaR5eFzgX3ebxm%2BHN7owoyTIl%2F1lA%3D&reserved=0</a><br>
<br>
Comparing the two versions with and without the requirement of ADC/TDC<br>
matching for the hodoscope, I only see a loss of events below 1% in<br>
almost the full energy range. There is only one 100MeV-wide bin near<br>
9.1GeV, where the loss is about 15%. Could this be a sign of a broken or<br>
miscalibrated TDC?<br>
<br>
For the 2 runs before the adjustment of the microscope thresholds, the<br>
loss of rho events with tagged beam photons is close to 30% as expected.<br>
Once the thresholds were fixed in run 71724, the loss goes down to 10%<br>
in the higher-energy part of the microscope, while it completely<br>
recovers for the low-energy part. What is the reason for this<br>
discrepancy? New vs old fibers?<br>
<br>
Best regards,<br>
<br>
Alex<br>
<br>
</blockquote></div>